A beam of 541nm light is incident on a diffraction grating that has 400grooves/mm. Determine the angle of the second-order ray.


1
Expert's answer
2021-03-03T11:30:37-0500

Answer

Grid spacing

"d=\frac{10^{-3}}{400}=2.50\times10^{-6}m"


And by equation

"dsin\theta=m\lambda"


"\theta=sin^{-1}(\frac{m\lambda}{d})"


"\theta=sin^{-1}(\frac{(2)(541\times10^{-9})}{250\times10^{-6})}) \\\theta=25.6°"